OpenStudy (anonymous):

I honestly don't understand how to solve linear word problems... Can someone help me solve one?

OpenStudy (anonymous):

I honestly don't understand how to solve linear word problems.. Here's the quesion: 1a. ) The Peanut gallery made a profit of $16,450 in 2002, and a profit of $21,340 in 2005. Assume that the profit followed a linear pattern. Write an equation that gives the profit, y, in terms of the year, x. Let x=0 correspond to 2000. Equation: 1b.) What was the profit in 2008 1c.) In what year would the profit be $31, 120 in the trend continued? Please show explanations & work. I really..don't understand this. (I obviously know the basics- solving equations, finding slope from an equation, y=mx + b, etc.)

OpenStudy (anonymous):

ok, so you know the generic linear equation y=mx+b, and you also know that y=profit, x = year (where x=0 corresponds to year 2000), also, you are given enough data to set up two linear equation of the form y=mx+b. can you try to set up two equations?

OpenStudy (anonymous):

if you need farther help, ask :D

OpenStudy (anonymous):

Um..I don't know..

OpenStudy (anonymous):

You're supposed to find one equation of this problem. That can be done, by using more than one equations I think..

OpenStudy (anonymous):

alright, so in the year 2002 (x=2), the profit is y = 16450, then that means: A: 16450 = 2m + b can you set up equation B?

OpenStudy (anonymous):

Equation "B" as in finding the slope?

OpenStudy (anonymous):

i mean another linear equation which i denote as "B", similar to equation "A" that I've set up above

OpenStudy (anonymous):

Okay. I need clarification. My Algebra teacher solved this problem like this: 

OpenStudy (anonymous):

Can you explain what she (teacher) did there?

OpenStudy (anonymous):

the equation is correct. what your teacher did is solve a system of two linear equations, which are the equations A which as shown earlier, and equation B, which has a very similar setup as A. to match what your teacher did, let A and B be two linear equations as shown below: A: y = mx + b B: Y = mX + b then, to solve for m: m = (Y-y)/(X-x)

OpenStudy (anonymous):

after solving for m, plug m back into either A or B to solve for b, and you get your equation :D

OpenStudy (anonymous):

What is equation "A" & "B"?

OpenStudy (anonymous):

equation "A" was shown earlier, remember? and "B" has a similar set up using the data given

OpenStudy (anonymous):

"alright, so in the year 2002 (x=2), the profit is y = 16450, then that means: A: 16450 = 2m + b can you set up equation B?"

OpenStudy (anonymous):

Oh. Yeah. Of course. Could you solve 1b for me?

OpenStudy (anonymous):

i can help you on 1b, but can you show me your 1a first :D

OpenStudy (anonymous):

Okay. y = 1630(2002) +13190

OpenStudy (anonymous):

^ is B

OpenStudy (anonymous):

for B the year would be 2005, so x = ?

OpenStudy (anonymous):

2002 is for A, which makes x = 2, what about B?

OpenStudy (anonymous):

Yeah..I think I did something wrong with this equation

OpenStudy (anonymous):

remember, x=0 mean 2000, that means x=1 is 2001, x=2 is 2002, x=3 is 2003 and so on

OpenStudy (anonymous):

y = 1630(8) + 13190

OpenStudy (anonymous):

thats for 1b, which means y = ?

OpenStudy (anonymous):

y= 13040 + 13190 y = 26230

OpenStudy (anonymous):

26230 = profit.

OpenStudy (anonymous):

yep you got it :D

OpenStudy (anonymous):

Although...?

OpenStudy (anonymous):

$26230

OpenStudy (anonymous):

Where did '5- 2" from the upper equation used to solve 1a come from?

OpenStudy (anonymous):

that is from "X-x", where X = 5 for 2005, and x = 2 for 2002

OpenStudy (anonymous):

Why did they need to be subtracted from each other?

OpenStudy (anonymous):

Why did "5-2" have to happen? Why is that relevant?

OpenStudy (anonymous):

What is B - A?

OpenStudy (anonymous):

What..?

OpenStudy (anonymous):

What is equation B minus equation A?

OpenStudy (anonymous):

2005 - 2002? --> Used 5 from 2005, and "2" from 2002?

OpenStudy (anonymous):

I don't know. That's why I'm asking you.

OpenStudy (anonymous):

What is equation A and equation B? you should know this by now.

OpenStudy (anonymous):

but, if you just want to understand the X-x part, then you have to understand what the slope m of an linear equation means

OpenStudy (anonymous):

it means the rate of change of y with respect to x

OpenStudy (anonymous):

Yes I know. But..

OpenStudy (anonymous):

I'm very frustrated. I was just wondering if you could solve the equation, (and just the equation) and give me step-by-step explanations.

OpenStudy (anonymous):

Your teaching methods..aren't helping me...that much. No offense.......

OpenStudy (anonymous):

To solve for the slope m, just compute m = (Y-y)/(X-x) where Y and X are from one linear equation, while y and x are from the other. then substitute m back into anyone of those two equations, to solve for b, and you get the equation that you want.